OpenStudy (anonymous):

exponent3(sqrt)1/8x exponent 12 y exponent 21 how do i solve??

OpenStudy (anonymous):

hey baby help me lol i need to solve

OpenStudy (anonymous):

ok baby i will

OpenStudy (anonymous):

ok baby i will

OpenStudy (anonymous):

ok baby i will

OpenStudy (anonymous):

ok like i dont undersand the steps to solve it

OpenStudy (anonymous):

(9 y Exponent^3)/sqrt(x)

OpenStudy (anonymous):

how do you solve 2-1 the -1 is an exponent *4 exponent 2-2

OpenStudy (anonymous):

hey baby you got a good voice

OpenStudy (anonymous):

0.9 L (liters)

OpenStudy (anonymous):

that is for the other problem

OpenStudy (anonymous):

thanks what is the diffrence between a centimeter and a centi gram??

OpenStudy (anonymous):

A centigram is a measure of mass, with dimensions and A centimeter is a unit of measurement in the metric system that is equal to one hundredth of a meter

OpenStudy (anonymous):

and how do i solve \[2^{-1} \times4^{2}-2\]

OpenStudy (anonymous):

what happen to the video chat and thanks for the medal

OpenStudy (anonymous):

i got in trouble ok i still dont understand about square roots and ur welcome and how do i solve that equation sorry im like multi tasking from algebra1 to biology to world history to health uugh fml

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

i like it

OpenStudy (anonymous):

= 30

OpenStudy (anonymous):

ok so explain the point of exponents and square roots combine like \[\sqrt{2^{4}}-3\times1\]

OpenStudy (anonymous):

idfk how to do this uugh algebra makes me wanna cuss ..lolz

OpenStudy (anonymous):

i am doing algegra 2

OpenStudy (anonymous):

at this rate im never going to pass acts...........algebra 2.................. :'(

OpenStudy (anonymous):

what is acts

OpenStudy (anonymous):

a c t s

OpenStudy (anonymous):

dont worry i am going to help you as best as i can

OpenStudy (anonymous):

not helping lolz

OpenStudy (anonymous):

=4

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i wish i was there

OpenStudy (anonymous):

me to wtf is a scientific notation

OpenStudy (anonymous):

can u hear me??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

know i got a school problem

OpenStudy (anonymous):

ok well what is it im good at everything but algebra

OpenStudy (anonymous):

first my school messed up my classes for the second time

OpenStudy (anonymous):

uugh ok nevermind lolz i cant help with that lolz

OpenStudy (anonymous):

thanks for being so sweet and nice to me

OpenStudy (anonymous):

no prob im ur girl its wat i do if its says evalute what does it mean

OpenStudy (anonymous):

love u gotta go bye ba and my stuff is courtneyfayby

OpenStudy (anonymous):

bye baby i love you

OpenStudy (anonymous):

hey baby what is a scientific notation

OpenStudy (anonymous):

baby?? what is a scientific notation

OpenStudy (anonymous):

Scientific notation is a way to "easily" or "conveniently" write very large or very small numbers

OpenStudy (anonymous):

what is radiocarbon dating??

OpenStudy (anonymous):

u still didnt answer my question .......

OpenStudy (anonymous):

baby i am sorry i was comparing you to her a chemical analysis used to determine the age of organic materials based on their content of the radioisotope carbon-14

OpenStudy (anonymous):

ok english please lolz im naturally blond and thats ok but it upset me and kinda hurt my feelings

OpenStudy (anonymous):

im really really really sorry

OpenStudy (anonymous):

i love you lol stop being a jerk today

OpenStudy (anonymous):

promise

OpenStudy (anonymous):

i promise to not be a jerk today

OpenStudy (anonymous):

i love you to i miss u and been thinking about all day

OpenStudy (anonymous):

k and awe so why are u being meun u ok ?? upset ??

OpenStudy (anonymous):

mean

OpenStudy (anonymous):

i am a little upset but i not mad at u sorry if i was acting that way

OpenStudy (anonymous):

to u

OpenStudy (anonymous):

do you have anymore questions

OpenStudy (anonymous):

awe whats bothering u honey and yea i think im gonna drop out and get my ged will you be mad if i do

OpenStudy (anonymous):

i wont be man at u but are you going to college and i will support you no matter what you decide to do i hope you feel better and i wish i was there to make you feel better

OpenStudy (anonymous):

and its something that happened in school

OpenStudy (anonymous):

<3 awe ok and yea i want to go to colledge but idk i think im gonna join the marines and tell me baby and how do you convert 45700 into a scientific notation

OpenStudy (anonymous):

= 45700.00 me + you = forever

OpenStudy (anonymous):

awe baby so what happened at school that has you so upset??

OpenStudy (anonymous):

everyday for the past couple of years the people that i know just make fun

OpenStudy (anonymous):

baby what do you like to do for fun

OpenStudy (anonymous):

story of my life lolz and i sing dance and im into football and others sports

OpenStudy (anonymous):

i want to hear you sing sometime 3-4x=7-5x =.x=7.

OpenStudy (anonymous):

i just want to leave fl and be with u

OpenStudy (anonymous):

3(4-x)-22=x+48 = x = -29/2

OpenStudy (anonymous):

awe ok baby so my dad said if i let him do a background check on u hes ok with us being together beacuse its a long distance relationship and we met online he wants to make sure ur you ya know do u have anything to hide or anything to tell me i want to kno before he does it u understand why id let him do it right?? your not mad at me are you i mean i want his permission to be together i love you and i wanna be with you forever ...anything to hide??

OpenStudy (anonymous):

baby??...

OpenStudy (anonymous):

(x-3)-(-2x+5=2x-7-3x) = -(5-2 x = -x-7)+x-3,, and ok he can do that and i have never got in trouble with the law i dont drink or do drugs i not at you my mom is the same the only thing i ever did was talk like $5 from my mom i will never

OpenStudy (anonymous):

awe i love you baby thankyou

OpenStudy (anonymous):

and is this only your dad or is it your dad and mom that wants to do the background check i wanna be with you forever to and i dont party never been in a fight i love u to ur welcome

OpenStudy (anonymous):

(5x-2)-(6-2x)=-x-9+5x = 7 x-8 = 4 x-9 x = -1/3.

OpenStudy (anonymous):

3 1/2x+17=2 3/4x+12+1/2x = (7 x)/2+17 = (11 x)/4+1/(2 x)+12 x = 1/3 (-10-sqrt(106))

OpenStudy (anonymous):

baby u there

OpenStudy (anonymous):

10x-7=9-4x = x = 8/7.

OpenStudy (anonymous):

awe baby and my dad and partially my mom im glad your letting me do it does it bother you at all??

OpenStudy (anonymous):

no it does not and even dough we meet online i will never hurt you and i will allways take care of you

OpenStudy (anonymous):

is there anything else your parents want to ask

OpenStudy (anonymous):

awe baby btw i gotta log off in 15 minutes and idk add my dad on face book eric fay his profile pic is of me and him last october talk to him get to know him k

OpenStudy (anonymous):

ok baby

OpenStudy (anonymous):

we need to set a wedding date so ik how much time i have to pick out a dress and do invites and stuff

OpenStudy (anonymous):

does your dad have a yellow hat on in his proflie

OpenStudy (anonymous):

no ... im with him his hair is down and im in a green shirt

OpenStudy (anonymous):

yea we do do you know how old you have to be to get married baby

OpenStudy (anonymous):

at least 17 here

OpenStudy (anonymous):

18 here

OpenStudy (anonymous):

14 in vegas ...lolz i want a fall wedding what do you think

OpenStudy (anonymous):

maybe october 18 2013 i gotta go goodnight baby

OpenStudy (anonymous):

i want you to have the wedding of your dreams did my friend/neighbor tom add you hes a good person i have known him pretty much my whole life